Let ABCD is a square whose two vertices A and D lie on positive x-axis and positive y-axis respectively.If the co-ordinate of C is(15,19),then the co-ordinate of B are:
1. (19,4)
2.(4,19)
3.(19,15)
4.(19,11)

Grade:12th Pass

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans: (19, 4)
Sol:
Let the coordinates of A, B & D be:
A(a, 0), B(c, d), D(0, b)
Equate the sides of the square:
a^{2}+b^{2}=15^{2}+(19-b)^{2}=(a-c)^{2}+d^{2}=(15-c)^{2}+(19-d)^{2}Equating diagonals of square:
(15-a)^{2}+(19)^{2}=c^{2}+(d-b)^{2}
After solving all the equations, you would find
a = 4
b = 15
c = 19
d = 4
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty